### v2

# Towards Underwater Camouflaged Object Tracking: Benchmark and Baselines

Abstract
Over the past decade, significant progress has been made in visual object tracking, largely due to the availability of large-scale datasets. However, existing tracking datasets are primarily focused on open-air scenarios, which greatly limits the development of object tracking in underwater environments. To bridge this gap, we take a step forward by proposing the first large-scale multimodal underwater camouflaged object tracking dataset, namely UW-COT220. Based on the proposed dataset, this paper first comprehensively evaluates current advanced visual object tracking methods and SAM- and SAM2-based trackers in challenging underwater environments. Our findings highlight the improvements of SAM2 over SAM, demonstrating its enhanced ability to handle the complexities of underwater camouflaged objects. Furthermore, we propose a novel vision-language tracking framework called VL-SAM2, based on the video foundation model SAM2. Experimental results demonstrate that our VL-SAM2 achieves state-of-the-art performance on the UW-COT220 dataset. The dataset and codes can be accessible at https://github.com/983632847/Awesome-Multimodal-Object-Tracking.
1 Introduction
Visual object tracking (VOT) involves continuously locating a target object within a video sequence and has applications in autonomous vehicles xie2024diffusiontrack ; wang2024elysium ; wang2024event , surveillance zhang2024awesome ; cai2024hiptrack , and robotics zhang2022webuav ; zhang2023all . Its significance lies in enabling machines to perceive and interpret dynamic environments, supporting tasks like motion analysis and decision-making li2024visual ; hu2024multi ; peng2024vasttrack . Although significant progress has been made in terrestrial and open-air scenarios huang2019got ; fan2021lasot , tracking in underwater environments remains challenging due to factors like visual camouflage, light scattering, and low contrast, which limit the effectiveness of conventional algorithms zhang2024webuot ; alawode2023improving ; alawode2022utb180 . Consequently, there is a pressing need for specialized datasets and methods to tackle the complexities of underwater tracking, especially when objects blend with their surroundings, known as camouflaged objects. However, despite its importance, underwater camouflaged object tracking remains an unexplored field.
Current VOT methods can be broadly categorized into several types. Traditional correlation filter-based methods zhang2019robust ; ge2020cascaded ; ge2019distilling ; zhang2020accurate utilize efficient convolution operations to track objects, whereas Siamese-based methods bertinetto2016fully ; li2019siamrpn++ ; li2018high employ dual-stream networks to learn a similarity function between the target object and the search region. Transformer-based methods, such as OSTrack ye2022joint , SeqTrack chen2023seqtrack , and ARTrack wei2023autoregressive , leverage the attention mechanism to model complex dependencies in sequences, resulting in enhanced tracking precision. In recent years, Mamba-based approaches like Mamba-FETrack huang2024mamba , MambaVT lai2024mambavt , and MambaTrack zhang2024mambatrack have emerged, offering robust performance through feature enhancement and advanced attention mechanisms. In addition to these methods, SAM-based models zhang2023comprehensive ; zhang2024segment have also gained traction. These models, such as SAM-DA fu2023sam , Tracking Anything yang2023track , SAM-Track cheng2023segment , and the latest SAMURAI-L yang2024samurai , focus on the precise segmentation of objects across frames, making them particularly suitable for challenging environments like underwater scenarios where object boundaries are difficult to discern.
Figure 1: Some examples come from our UW-COT220 dataset. We provide bounding box and language description annotations for a camouflaged underwater target object in each video sequence.
Table 1: Comparison of the constructed UW-COT220 with existing camouflaged object tracking dataset (i.e., COTD guo2024cotd ) and video camouflaged object detection datasets (i.e., CAD bideau2016s and MoCA-Mask cheng2022implicit ). UW-COT220 is the first multimodal underwater camouflaged object tracking dataset.
  Dataset Year Videos Classes
Min
frame
Mean
frame
Max
frame
Total
frames
Labeled
frames
Annotation Link
Language
descriptions
Underwater
CAD bideau2016s 2016 9 6 30 93 218 836 191 Mask URL
MoCA-Mask cheng2022implicit 2022 87 45 23 264 1,296 23 K 23 K Mask URL
COTD guo2024cotd 2024 200 20 - 400 - 80 K 80 K BBox URL
UW-COT220 (Ours) 2024 220 96 10 722 7,448 159 K 159 K BBox+Mask URL
Following the latest large-scale underwater object tracking dataset (i.e., WebUOT-1M zhang2024webuot ), we take a step forward and construct the first multimodal underwater camouflaged object tracking dataset, UW-COT220 (see Fig. 1 and Tab. 1), as a benchmark for evaluating and developing advanced underwater camouflaged object tracking methods. The dataset consists of 220 underwater video sequences, spanning 96 categories, with approximately 159,000 frames. We annotate each video with a language description to promote the exploration of multimodal underwater camouflaged object tracking. As a single-object visual tracking task, we provide bounding box annotations for the camouflaged underwater objects in each frame. Additionally, using manually annotated bounding boxes as prompts, we generate pseudo mask annotations for the camouflaged underwater objects by leveraging advanced interactive segmentation models (e.g., HQ-SAM ke2024segment and SAM2 ravi2024sam ). We conduct pioneering experiments on the constructed dataset to evaluate the performance of state-of-the-art visual object tracking methods and the latest advancements in image and video segmentation. By releasing UW-COT220, we aim to inspire interest in exploring various underwater vision tasks and to foster the development of advanced methods for underwater camouflaged object tracking.
2 UW-COT220 Dataset
Our goal is to construct a large-scale underwater camouflaged object tracking dataset that involves a rich variety of categories and various real underwater scenes for evaluating and developing general underwater camouflaged object tracking methods. To achieve this, we collect underwater videos from video-sharing platforms (e.g., YouTube [^1] https://www.youtube.com/) and existing tracking datasets (e.g., WebUOT-1M zhang2024webuot and VastTrack peng2024vasttrack ), and we filter out 220 videos that contain underwater camouflaged objects. Following existing datasets huang2019got ; fan2019lasot ; zhang2022webuav ; peng2024vasttrack ; zhang2024webuot , we provide a bounding box annotation for the camouflaged object in each frame, represented as $[x,y,w,h]$, where $x$ and $y$ denote the coordinates of the top-left corner of the object, and $w$ and $h$ indicate the width and height of the object. In each video, we annotate an English sentence describing the color, behavior, attributes, and surroundings of the target. Furthermore, to enhance the accuracy of object representation, we use segmentation foundation models (e.g., HQ-SAM ke2024segment and SAM2 ravi2024sam ) to generate pseudo mask annotations for the camouflaged objects. Some examples and detailed statistics of the UW-COT220 dataset are provided in Fig. 1 and Tab. 1. To the best of our knowledge, UW-COT220 is the first large-scale benchmark for underwater camouflaged object tracking, featuring a diverse set of 96 categories and various challenging underwater scenes.
Figure 2: Overview of the proposed vision-language tracking framework VL-SAM2. Our VL-SAM2 retains the core architecture of SAM2 and freezes the weights of the image encoder. The language prompt is encoded by the language encoder and then injected into the mask decoder. Additionally, a training-free MATP module is integrated to enhance tracking robustness during inference.
3 Proposed Method
In this section, we introduce VL-SAM2, a new vision-language tracking framework for multimodal underwater camouflaged object tracking, as shown in Fig. 2. The proposed VL-SAM2 is based on the latest video foundation model, SAM2, and extends its capabilities to challenging underwater scenarios. VL-SAM2 consists of three main components: a visual branch (i.e., SAM2), a language branch, and a training-free motion-aware target prediction (MATP) zhang2024webuot module.
The input to VL-SAM2 consists of video frames and a language description. First, we use an image encoder and a language encoder to extract features from both modalities. In the first frame, the initial bounding box (BBox) is encoded by the prompt encoder to obtain the visual embedding. The language embedding is processed by a linear projection to match the dimension of the visual embedding. Then, we concatenate the visual and language embeddings as sparse embeddings and feed them into the mask decoder. In subsequent frames, we adopt a video propagation manner similar to SAM2 to obtain the prediction result for each frame. The MATP, based on Kalman filtering kalman1960new , is used to alleviate the model drift caused by similar distractors around the true target. This module was proposed in zhang2024webuot . Inspired by yang2024samurai , we further apply MATP to update the memory of SAM2. We strongly recommend readers refer to zhang2024webuot and yang2024samurai for more details.
We use Hiera-L ryali2023hiera as the image encoder and CLIP ViT-B/32 radford2021learning as the language encoder. During training, we freeze the parameters of both the image encoder and the language encoder, and fine-tune the other modules. This allows us to train the proposed VL-SAM2 on an Ubuntu server with 8 A6000 GPUs. The current version has been trained only on the Ref-Youtube-VOS seo2020urvos dataset and then evaluated in a zero-shot manner on the UW-COT220 dataset.
Figure 3: Overall performance on the UW-COT220 dataset using AUC, Pre, cAUC, and nPre scores.
Table 2: Impact of using different points (within the initial target box) as the point prompt for SAM2 ravi2024sam . We report AUC, nPre, Pre, cAUC, and mACC scores on the UW-COT220 dataset.
  Method Point type AUC (%) Pre (%) nPre (%) cAUC (%) mACC (%)
SAM2-tiny Center point 53.1 53.7 60.1 52.2 53.4
Random point 37.7 37.2 41.6 36.4 37.3
SAM2-small Center point 52.8 53.6 60.2 52.0 53.1
Random point 36.3 35.3 41.0 35.2 36.0
SAM2-base plus Center point 55.0 56.0 62.2 54.2 55.4
Random point 40.5 40.4 45.8 39.5 40.2
SAM2-large Center point 60.6 63.5 67.3 60.0 61.3
Random point 44.6 45.5 49.3 43.5 44.6
Table 3: Comparison of different models for SAM2 ravi2024sam on the UW-COT220 dataset.
  Method Size (M) Speed (FPS) AUC (%) Pre (%) nPre (%) cAUC (%) mACC (%)
SAM2-tiny 38.9 47.2 53.1 53.7 60.1 52.2 53.4
SAM2-small 46.0 43.3 52.8 53.6 60.2 52.0 53.1
SAM2-base plus 80.8 34.8 55.0 56.0 62.2 54.2 55.4
SAM2-large 224.4 24.2 60.6 63.5 67.3 60.0 61.3
4 Experimental Results
Main Results. We evaluate SAM-based tracking methods (i.e., SAM-DA fu2023sam , and Tracking Anything yang2023track ), SAM2, and several current SOTA VOT methods (i.e., OSTrack ye2022joint , SeqTrack chen2023seqtrack , ARTrack wei2023autoregressive , and SAMURAI-L yang2024samurai ) on the proposed UW-COT220 dataset. The results are shown in Fig. 3. Our observations are as follows: 1) Our VL-SAM2 achieves the best performance, surpassing the SOTA trackers, including SAMURAI-L yang2024samurai and SAM2. 2) The top three trackers (VL-SAM2, SAMURAI-L, and SAM2) are all SAM2-based, highlighting the promising trend of using foundation models for tracking. 3) SAM2-based methods outperform SAM-based trackers (SAM-DA and Tracking-Anything) on UW-COT220, which can be attributed to a series of improvements SAM2 introduces over SAM for video and image tasks, such as improving temporal consistency, robustness to occlusions, feature embeddings, computational efficiency, motion estimation accuracy, generalization to new domains, and integration of contextual information.
Impact of Prompt. We take SAM2 as our baseline tracker to explore the impact of different ways of point prompts (i.e., center point and random point within the initial target box). From Tab. 2, we find that using the center point as a prompt yields significantly better results than using random points. This implies that for the interactive segmentation model SAM2, the quality of the input prompt plays a crucial role in determining its performance. This also motivates us to explore the use of more flexible prompts, e.g., language descriptions, in the context of underwater camouflaged object tracking.
Impact of Model Size. The results in Tab. 3 demonstrate that larger model sizes generally lead to better performance, but the speed of the model decreases significantly. We also discovered an interesting phenomenon: when the number of model parameters is relatively small, a smaller model (e.g., SAM2-tiny) can even outperform a larger model (e.g., SAM2-small). We suspect this may be due to overfitting, or that small models (e.g., SAM2-small) are more sensitive to the quality of the training data ravi2024sam .
Impact of Language Encoder. In Tab. 4, we explore the impact of using different language encoders in the proposed VL-SAM2 framework. T5 (Text-to-Text Transfer Transformer) raffel2020exploring is a large-scale language model that models all natural language processing tasks as text-to-text problems, while CLIP radford2021learning is a contrastive language-image pre-training model. The results demonstrate that using CLIP as the language encoder outperforms T5. We hypothesize that this is due to CLIP’s text encoder being aligned with the visual encoder in the feature space over a vast amount of data, making it more suitable for the VOT task.
Table 4: Impact of different language encoders on the UW-COT220 dataset.
  Method Language Encoder AUC (%) Pre (%) nPre (%) cAUC (%) mACC (%)
VL-SAM2 T5 raffel2020exploring 54.6 54.8 59.3 53.9 55.1
VL-SAM2 CLIP radford2021learning 71.3 75.2 78.2 70.8 72.4
5 Conclusion
In this paper, we introduce UW-COT220, the first large-scale multimodal benchmark for underwater camouflaged object tracking, and demonstrate the superior performance of SAM2 over other tracking methods, showcasing its potential for enhancing underwater tracking technologies. To promote research on multimodal underwater camouflaged object tracking, we propose a simple yet effective vision-language tracking framework called VL-SAM2. In the future, we plan to expand the scale and modalities of our dataset, as well as explore various underwater vision tasks.